A player pays $10 to roll two six-sided dice. She wins $50 if she rolls a sum of 10 and $20 if she rolls a sum of 8. Given the probability distribution for the game (below), compute the expected value of the game to the nearest cent.
Answer #1

Solution:

P(Sum of 8) = 6/36 = 1/6

P(Sum of 10) = 2/36 = 1/18

P(Rest any sum) = 1 - (6 + 2)/36

= 1 - 8/36

= 1 - 2/9

= 7/9

Hence,

Expected value of the game

= $(50 - 2)*(1/6) + $(20- 2)*(1/18) - $10*(7/9)

= $ 1.222

Note: Please comment below if you have any doubts